Give a worked out solution, will rate if correct

An x-ray photon with wavelength 6.07 pm collides with a free electron initially at rest. What is the maximum possible kinetic energy acquired by the electron?
Answer #1

Maximum energy acquired by electron= energy of photon= hc/?

=6.62*10-34*3*108/6.07*10-12=3.275*10-14J=0.2044 MeV